# Peer Review Status: Submitted to Seismica

### v1

@@
Abstract
In theory, earthquake magnitudes follow an exponential distribution. In practice, however, earthquake catalogs report magnitudes with finite resolution, resulting in a discrete (geometric) distribution. To determine the lowest magnitude above which seismic events are completely recorded, the Lilliefors test is commonly applied. Because this test assumes continuous data, it is standard practice to add uniform noise to binned magnitudes prior to testing exponentiality.
This work shows analytically that uniform dithering cannot recover the exponential distribution from its geometric form. It instead returns a piecewise-constant residual lifetime distribution, whose deviation from the exponential model becomes detectable as catalog size or bin width increases. Numerical experiments confirm that this deviation yields an overestimation of the magnitude of completeness in large catalogs.
We therefore derive the exact noise distribution - a truncated exponential on the bin interval - that correctly restores the continuous exponential distribution over the whole magnitude range. Numerical tests show that this correction yields Lilliefors rejection rates consistent with the significance level for all bin widths and catalog sizes. The proposed solution eliminates a methodological bias in completeness estimation, which especially impacts high-resolution catalogs.
@@ 1 Introduction
The cumulative (survival) frequency-magnitude distribution of earthquakes is described by the Gutenberg–Richter (GR) law (gutenberg1944frequency), a decreasing exponential model valid for all magnitudes above a given threshold known as the magnitude of completeness. Below this threshold, the distribution deviates from exponential behavior due to the incomplete recording of low-magnitude events, which is influenced by the quality and configuration of the seismic network (Seismic Network Density Incompleteness, SNDI) and by transient increases in noise following large earthquakes (Short-Term Aftershock Incompleteness, STAI) (mignan2012theme, and references therein). Accurately estimating the magnitude of completeness - and thus identifying the range where the exponentiality assumption holds - is essential for any statistical analysis based on earthquake catalogs.
@@
The exponentiality of earthquake magnitudes should not be assumed a priori, but instead statistically verified. A common approach for testing the exponentiality of continuous data is the Lilliefors test (lilliefors1969kolmogorov; clauset2009power), a modification of the Kolmogorov–Smirnov (KS) test (massey1951kolmogorov) designed for cases where parameters are estimated from the data. However, magnitudes reported in real seismic catalogs are discrete rather than continuous, typically provided with a resolution of one decimal place. To address this, a standard practice is to add a random variable uniformly distributed in the range $\Delta m/2,+\Delta m/2]$ (where $\Delta m$ is the bin width), which transforms discrete magnitudes into continuous variables. It has been noted, however, that for $\Delta m>0.01$, this procedure may compromise the reliability of the Lilliefors test and bias the estimates of the magnitude of completeness (herrmann2021inconsistencies; spassiani2023real).
@@
This work was initially motivated by the observation that the Lilliefors test can return unexpectedly high estimates of the magnitude of completeness when applied to enhanced catalogs, i.e., high-resolution catalogs produced using template-matching (e.g. gibbons2006detection; shelly2007non; chamberlain2018eqcorrscan) or machine learning techniques (e.g. dokht2019seismic; zhu2019phasenet; mousavi2020earthquake). This may come as a surprise, given that such enhanced catalogs primarily increase the number of recorded events at lower magnitudes. However, the higher rejection rate observed could, in principle, reflect true deviations from exponentiality at low magnitudes caused, for example, by improper mixing of magnitude types (herrmann2021inconsistencies), or by the greater uncertainty associated with small magnitudes (pattonexploring). While these issues could affect routine catalogs as well, the substantially larger amount of data in enhanced catalogs - usually containing more than ten times the number of events found in routine catalogs - could explain the higher rejection rates of the Lilliefors test, making it more sensitive to even subtle deviations from the null hypothesis of exponentiality (increased statistical power). Addressing this problem is crucial, as it can strongly affect any statistical analysis performed on high-resolution catalogs relying on the proper estimation of the magnitude of completeness (herrmann2021inconsistencies; mancini2022use; stallone2025comparison).
@@
An alternative explanation, however, is that the higher rejection rate actually reflects a methodological flaw in how magnitude exponentiality is tested. To rule this out, we investigate whether the common procedure of dithering binned magnitudes with uniform noise, prior to applying the Lilliefors test, systematically introduces spurious deviations from exponentiality. We address this problem both analytically and numerically, and demonstrate that dithering binned magnitudes with uniformly distributed noise fails to correctly recover the exponential distribution. Then, we derive the appropriate distribution that the added random variable must follow to ensure a correct transformation from discrete to continuous exponentially distributed values. Finally, we validate the proposed method through numerical experiments.
@@ 1.1 Binned magnitudes
$m_{i}=m_{0}+\left(i-\frac{1}{2}\right)\Delta m,\hskip 18.49988pti=1,2,\dots$ (1)
@@
$p_{M}(m)=\beta e^{-\beta(m-m_{0})},\hskip 18.49988ptm\geq m_{0}.$ (2)
@@
Remark. We recall that the geometric distribution has the interpretation of the number of failures in a sequence of Bernoulli trials until the first success. If we consider a regime where the probability of success is very small, such that $p=\frac{\lambda}{N}$, in the large $N$ limit it can be proved that the geometric distribution converges to the exponential distribution with parameter $\lambda$ (e.g. feller1957introduction).
@@ 1.2 Uniformly dithered (binned) magnitudes
As discussed in Section 1.1, although theoretical magnitudes follow a continuous exponential distribution, in practice they are represented as binned values, consistent with the discrete resolution of real catalogs. When continuous data are required for statistical tests, such as using the Lilliefors test to estimate the magnitude of completeness, the binned magnitudes are commonly dithered by adding uniform noise. In fact, the distribution of the random variable (r.v.) $M=M_{i}+Y$, where $M_{i}$ is discrete while $Y$ is continuous, is effectively continuous. To prove this result, we recall that a generic random variable $M$ is continuous if and only $\mathbb{P}(M\in A)=0$ for all Borel measurable sets $A$ with zero Lebesgue measure. Now, if $M_{i}$ has PMF $p_{i}$ and $Y$ has PDF $f_{Y}(y)$, the distribution of $M=M_{i}+Y$ is given by
@@
$0\leq\mathbb{P}(M\in A)=\sum_{i}\mathbb{P}(Y\in A-m_{i}\cap M_{i}=m_{i})\leq\sum_{i}\mathbb{P}(Y\in A-m_{i})=0,$
where the last equality follows from the fact that the set $A-m_{i}$ is still Lebesgue null. Therefore, it follows that $\mathbb{P}(M\in A)=0$ and $M$ is a continuous r.v.
In our specific case, we have
@@
The resulting continuous distribution for $M$ can therefore be obtained in the following way.
@@
Since $f_{Y}(m-i)\neq 0$ only if $(m-i)\in(a,b)$, or equivalently $m\in(a+i,b+i)$, the support of $f_{M}(m)$ is the union of the intervals $(i+a,i+b)$, for $i\in\mathbb{N}^{+}$. Then
@@
where we used $q=1-e^{-\beta\Delta m}$. This is a piecewise constant, staircase PDF with values in the interval $(1+a,\infty)$, with decreasing steps. It precisely consists of the sum of rectangular bumps of width bab-a, centered at every $i\in\mathbb{N}^{+}$, with exponentially decaying height
$\frac{1-e^{-\beta\Delta m}}{b-a}\,e^{-\beta\Delta m\,(i-1)},\qquad m\in\left(a+i,b+i\right),\qquad i\in\mathbb{N}^{+}.$
@@
that is, the classical magnitude noise, uniformly distributed in $(-\frac{\Delta m}{2},\frac{\Delta m}{2})$, we get
@@
The sum above is non-zero only for the integer $i_{m}\in\mathbb{N}^{+}$ such that $m\in\left(i_{m}-\frac{\Delta m}{2},i_{m}+\frac{\Delta m}{2}\right)$, so the final expression for $f_{M}(m)$ is
$f_{M}(m)$ $=\begin{cases}\frac{1-e^{-\beta\Delta m}}{\Delta m}\,e^{-\beta\Delta m\,(i_{m}-1)},&\text{if}\quad m\in\left(i_{m}-\frac{\Delta m}{2},i_{m}+\frac{\Delta m}{2}\right),\quad i_{M}\in\mathbb{N}^{+}\\ 0&\text{otherwise},\end{cases}$ (10)
@@
We can conclude that, if $M_{i}$ is geometrically distributed with parameter $1-e^{-\beta\Delta m}$, while $Y$ is uniformly distributed in $(-\frac{\Delta m}{2},\frac{\Delta m}{2})$, the random variable $M=M_{i}+Y$ follows a residual lifetime distribution, whose PDF $f_{M}(m)$:
@@
Limiting case. To compute the limit, for $\Delta m\to 0$, of the function $f_{M}(m)$ defined in (10), we can consider the property $\lim_{x\to 0}\frac{s^{tx}-1}{x}=t\,\ln s$, which gives
$\lim_{\Delta m\to 0}\frac{1-e^{-\beta\Delta m}}{\Delta m}=\lim_{\Delta m\to 0}\beta\frac{e^{\beta\Delta m}-1}{\beta\Delta m}=\beta.$
Now, recalling that $m-\frac{\Delta m}{2}-m_{0}=\Delta m(i-1)$ (see Eqs. (1)-(4)), we can observe that the classical exponential PDF is recovered when the bin width $\Delta m$ approaches 0. In this limiting case, the exponential and uniform mean points coincide within each $\Delta m$-bin used to group continuous magnitudes sampled from an exponential distribution. This result also follows directly from the memoryless property of the exponential distribution (marzocchi2019how).
@@ 1.2.1 Finite magnitudes’ dataset (real case-studies)
The calculations presented above are based on an infinite, discrete set of magnitudes, i.e., $i=1,2,\dots$. However, this assumption does not hold in real applications, where earthquake catalogs are finite and may vary significantly in quality and size - from statistically limited small datasets to more recent high-resolution records. In theoretical terms, this implies considering a truncated geometric distribution for $M_{i}$. Specifically,
@@
where we used $q=1-e^{-\beta\Delta m}$. This is again a piecewise constant, continuous PDF of a residual lifetime distribution, consisting of $N$ rectangular bumps of width $\Delta m$, centered at every $i=1,\dots,N$, with exponentially decaying height. Specifically, in each interval,
@@
Figure 1: Numerical test showing how $\Delta m$ impacts the approximation quality of the exponential distribution (solid black line). The analytical solutions of the piecewise constant staircase PDF (left) and CDF (right) are calculated for the following values of $\Delta m$: 0.1, 0.2, 0.3, 0.4, 0.5. Since the maximum magnitude is fixed ($M_{max}=5.0$), here $N$ is a dependent variable: $N=(M_{max}-M_{0})/\Delta_{M}$, where $M_{0}=1.0$ is the minimum magnitude.
@@ 1.2.2 Numerical test: binned magnitudes dithered with uniform noise
To simulate binned magnitudes, we draw samples from the discretized GR law, i.e. the geometric distribution with parameter $p=1-e^{-\beta\Delta m}$. We generate synthetic catalogs of binned magnitudes while varying only the catalog size and keeping all other parameters fixed: bin width $\Delta m=0.1$, bb\,-value =1=1, and minimum magnitude $M_{\min}=1.0$. For each catalog, we estimate the magnitude of completeness $M_{c}$ using the Lilliefors test ($\alpha=0.1$) as implemented in the Python routine Mc-Lilliefors (marcus_herrmann_2020_4162497; herrmann2021inconsistencies).
Table 1 shows that the Lilliefors test correctly identifies $M_{c}$ - i.e., $M_{c}=M_{\min}=1.0$ - only for catalog sizes on the order of several thousand events. For larger catalogs, the estimated $M_{c}$ becomes systematically higher, with the error increasing as the catalog size increases.
Table 1: Estimated magnitude of completeness $M_{c}$ using the Lilliefors test for synthetic catalogs with $M_{min}=1.0$ and varying size ($\alpha=0.1$).
@@
To examine this issue more in detail, we set up a numerical experiment to assess how the rejection rate of the Lilliefors test varies with the bin width $\Delta m$ and catalog size. For various combinations of $\Delta m$ and size, we generate 1,000 synthetic catalogs of binned magnitudes. Each catalog is then dithered following the standard procedure, i.e. by adding uniformly distributed noise. Lastly, we apply the Lilliefors test. The rejection rate is defined as the proportion of simulations in which the null hypothesis of exponentially distributed magnitudes is rejected. The full procedure is described in detail in Algorithm 1. Results are shown in Table 2 and Figure 3.
Algorithm 1 Lilliefors rejection rate as function of bin width and catalog size - Uniform noise
[…]
3:$b=1$ ( $\beta=b\cdot\ln(10)$ )
[…]
6:for each bin width $\Delta m$ do
7:   for each catalog size $n_{\mathrm{events}}$ do
8:    $counter=0$
9:    for $j=1$ to $N_{\mathrm{SIM}}$ do $\triangleright$ Simulate one catalog
10:     for $i=1$ to $n_{\mathrm{events}}$ do $\triangleright$ Generate binned magnitudes
11:        Draw $k_{i}\in\{0,1,2,\dots\}$ with $\Pr(k_{i}=k)=q(1-q)^{k}$, $q=1-e^{-\beta\Delta m}$
12:        Set $M_{i}=M_{\min}+k_{i}\cdot\Delta m$
13:     end for
14:     Initialize empty list $P$ $\triangleright$ store $p$-values
15:     for $k=1$ to $N_{\mathrm{NOISE}}$ do $\triangleright$ Dither magnitudes
16:        Draw noise $\varepsilon_{i}\sim\mathcal{U}(-\Delta m/2,\Delta m/2)$ for $i=1,\dots,n_{\mathrm{events}}$
17:        $M_{i}^{\mathrm{dithered}}=M_{i}+\varepsilon_{i}$
18:        Shift magnitudes: $M_{i}^{\mathrm{shifted}}=M_{i}^{\mathrm{dithered}}-(M_{\min}-\Delta m/2)$
19:        Apply Lilliefors test: $p_{k}=\mathrm{Lilliefors}(M^{\mathrm{shifted}},\mathrm{dist}=\mathrm{exp})$
20:        Append $p_{k}$ to $P$
21:     end for
22:     Compute mean p-value: $\bar{p}=\mathrm{mean}(P)$
23:     if $\bar{p}<\alpha$ then $counter\leftarrow counter+1$
24:     end if
25:    end for
26:    Rejection rate: $R=100\cdot\frac{counter}{N_{\mathrm{SIM}}}$
27:   end for
28:end for
Table 2: Lilliefors test rejection rates (%) for $\alpha$ = 0.1, based on 1,000 simulations with 100 noise realizations each, shown as a function of bin width ($\Delta m$) and catalog size.
@@
0.5 17 100 100 100 100
0.4 6 100 100 100 100
0.3 4 93 100 100 100
0.2 5 14 100 100 100
0.1 7 6 21 100 100
The rejection rate increases with both catalog size and bin width $\Delta m$. Specifically, for a fixed catalog size, the rejection rate tends to increase as $\Delta m$ increases. With the standard bin width ($\Delta m=0.1$), the exponential model is rejected 100% of the time for size of order $10^{5}$, which corresponds to the lower range of modern high‐resolution datasets (e.g. ross2019searching; tan2021machine).
@@
    The power of the Lilliefors test to detect the deviation increases with catalog size and bin width.
The apparently correct estimation of $M_{c}=M_{\min}$ (Table 1) and the low rejection rates at low-to-intermediate catalog sizes (Table 2) is an artefact caused by two opposing effects compensating each other, i.e. smoothing by dithering and the statistical power of the Lilliefors test. For catalog sizes of a few thousand events, the uniform noise smooths the geometric distribution sufficiently well that the deviations from the exponential model are too small for the Lilliefors test to detect, so the null hypothesis is not rejected. As the catalog size increases, however, the test gains statistical power and begins to detect the systematic deviations introduced by the uniform dithering, causing a growing number of rejections and an increase in the estimated $M_{c}$.
Figure 3: Lilliefors test rejection rates (%) for $\alpha$ = 0.1, computed from 1,000 simulations, each incorporating 100 noise realizations, across varying bin widths ($\Delta m$) and catalog sizes. See Table 2.

### v2

# Correcting exponentiality test for binned earthquake magnitudes

@@
Abstract Above the magnitude of completeness - the minimum threshold for which a 100% detection rate is assumed - earthquake magnitudes are typically modeled as a continuous exponential distribution. In practice, however, earthquake catalogs report magnitudes with finite resolution, resulting in a discrete (geometric) distribution. To determine the magnitude of completeness, the Lilliefors test is commonly applied. Because this test assumes continuous data, it is standard practice to add uniform noise to binned magnitudes prior to testing exponentiality. Here we show analytically that uniform dithering does not recover the underlying continuous exponential distribution from its discretized (geometric) form. It instead returns a piecewise-constant residual lifetime distribution, whose deviation from the exponential model becomes detectable as catalog size or bin width increases. Through numerical experiments, we demonstrate that this deviation yields a systematic overestimation of the magnitude of completeness, with biases exceeding one magnitude unit in large, high-resolution catalogs. We derive the exact noise distribution - a truncated exponential within each magnitude bin - that correctly restores the continuous exponential distribution over the whole magnitude range. Numerical tests show that this correction yields Lilliefors rejection probabilities that are consistent with the significance level across a wide range of bin widths and catalog sizes. Although illustrated for the Lilliefors test, the identified bias and the proposed correction are independent of the specific statistical test and apply generally to exponentiality testing of discretized magnitude data. {summary}Non-technical summary Earthquake magnitudes are commonly described by a continuous exponential distribution within the range where the Gutenberg-Richter law applies. However, reported magnitudes are discretized (binned), following a geometric distribution. The Lilliefors test is often used to estimate the magnitude of completeness, and common practice is to add uniform noise to discretized magnitudes so the test can be applied. This study shows that uniform noise does not recover the underlying exponential distribution and instead introduces a bias that becomes detectable for large catalogs or coarse binning, leading to overestimation of the magnitude of completeness. We derive the correct noise that must be added, i.e., a truncated exponential within each bin, and show that it restores the continuous exponential distribution.
@@ 1 Introduction
Within the magnitude range above the completeness threshold where self-similarity approximately holds, the complementary cumulative frequency-magnitude distribution of earthquakes is described by the Gutenberg–Richter (GR) law (gutenberg1944frequency), a decreasing exponential model. Below this threshold, the distribution deviates from exponential behavior due to the incomplete recording of low-magnitude events, which is influenced by the quality and configuration of the seismic network (Seismic Network Density Incompleteness, SNDI) and by transient increases in noise following large earthquakes (Short-Term Aftershock Incompleteness, STAI) (mignan2012theme, and references therein). Accurately estimating the magnitude of completeness - and thus identifying the range where the exponentiality assumption holds - is essential for any statistical analysis based on earthquake catalogs.
@@
The exponentiality of earthquake magnitudes should not be assumed a priori, but instead statistically verified. A common approach for testing the exponentiality of continuous data is the Lilliefors test (lilliefors1969kolmogorov; clauset2009power). Originally proposed to assess normality (lilliefors1967), the test was later adapted to exponential distributions. It is a modification of the Kolmogorov–Smirnov (KS) test (massey1951kolmogorov), specifically designed for situations in which the distribution parameters are estimated from the data and the observations are continuous. However, magnitudes reported in real seismic catalogs are discrete rather than continuous, typically provided with a resolution of one decimal place. To address this, a standard practice is to add a random variable uniformly distributed in the range $\Delta m/2,+\Delta m/2]$ (where $\Delta m$ is the bin width), which transforms discrete magnitudes into continuous variables. Previous studies have shown that, for $\Delta m>0.01$, this procedure may compromise the reliability of the Lilliefors test and bias the estimates of the magnitude of completeness (herrmann2021inconsistencies; spassiani2023real).
@@
This work was initially motivated by the observation that the Lilliefors test can return unexpectedly high estimates of the magnitude of completeness when applied to enhanced catalogs, i.e., high-resolution catalogs produced using template-matching (e.g. gibbons2006detection; shelly2007non; chamberlain2018eqcorrscan) or machine learning techniques (e.g. dokht2019seismic; zhu2019phasenet; mousavi2020earthquake). This may come as a surprise, given that such enhanced catalogs primarily increase the number of recorded events at lower magnitudes. However, the higher rejection probability observed could, in principle, reflect true deviations from exponentiality at low magnitudes caused, for example, by improper mixing of magnitude types (herrmann2021inconsistencies), or by the greater uncertainty associated with small magnitudes (pattonexploring). While these issues could affect routine catalogs as well, the substantially larger amount of data in enhanced catalogs - usually containing more than ten times the number of events found in routine catalogs - could explain the higher rejection probabilities of the Lilliefors test, making it more sensitive to even subtle deviations from the null hypothesis of exponentiality (increased statistical power). Addressing this problem is crucial, as it can strongly affect any statistical analysis performed on high-resolution catalogs relying on the proper estimation of the magnitude of completeness (herrmann2021inconsistencies; mancini2022use; stallone2025comparison).
@@
An alternative explanation, however, is that the higher rejection probability actually reflects a methodological flaw in how magnitude exponentiality is tested. To rule this out, we investigate whether the common procedure of dithering binned magnitudes with uniform noise, prior to applying the Lilliefors test, systematically introduces spurious deviations from exponentiality. We address this problem both analytically and numerically, and demonstrate that dithering binned magnitudes with uniformly distributed noise fails to correctly recover the exponential distribution. Then, we derive the appropriate distribution that the added random variable must follow to ensure a correct transformation from discrete to continuous exponentially distributed values. Finally, we validate the proposed method through numerical experiments.
@@ 1.1 Binned magnitudes
$m_{i}=m_{0}+\left(i-\frac{1}{2}\right)\Delta m,\qquad i=1,2,\dots$ (1)
@@
$p_{M}(m)=\beta e^{-\beta(m-m_{0})},\qquad m\geq m_{0}.$ (2)
@@ 1.2 Uniformly dithered (binned) magnitudes
As discussed in Section 1.1, although theoretical magnitudes follow a continuous exponential distribution, in practice they are represented as binned values, consistent with the discrete resolution of real catalogs. When continuous data are required for statistical tests, such as using the Lilliefors test to estimate the magnitude of completeness, the binned magnitudes are commonly dithered by adding uniform noise. In fact, the distribution of the random variable (r.v.) $M=M_{i}+Y$, where $M_{i}$ is discrete while $Y$ is continuous, is effectively continuous. To prove this result, we recall that a generic random variable $M$ is continuous if and only $P(M\in A)=0$ for all Borel measurable sets $A$ with zero Lebesgue measure. Now, if $M_{i}$ has PMF $p_{i}$ and $Y$ has PDF $f_{Y}(y)$, the distribution of $M=M_{i}+Y$ is given by
@@
$0\leq P(M\in A)=\sum_{i}P(Y\in A-m_{i}\cap M_{i}=m_{i})\leq\sum_{i}P(Y\in A-m_{i})=0,$
where the last equality follows from the fact that the set $A-m_{i}$ is still Lebesgue null. Therefore, it follows that $P(M\in A)=0$ and $M$ is a continuous r.v.
In our specific case, using the magnitude class representation (1) by tinti:1987, we have
@@
The resulting continuous distribution for $M$ can therefore be obtained in the following way:
@@
Since $f_{Y}(m-i)\neq 0$ only if $(m-i)\in(a,b)$, or equivalently $m\in(a+i,b+i)$, the support of $f_{M}(m)$ is the union of the intervals $(i+a,i+b)$, for $i\in N^{+}$. Then
@@
where we used $q=1-e^{-\beta\Delta m}$. This is a piecewise constant, staircase PDF with values in the interval $(1+a,\infty)$, with decreasing steps. It precisely consists of the sum of rectangular bumps of width bab-a, centered at every $i\in N^{+}$, with exponentially decaying height
$\frac{1-e^{-\beta\Delta m}}{b-a}\,e^{-\beta\Delta m\,(i-1)},\quad m\in\left(a+i,b+i\right),\quad i\in N^{+}.$
@@
that is, the classical magnitude noise uniformly distributed in $(-\frac{\Delta m}{2},\frac{\Delta m}{2})$, we get
@@
The sum above is non-zero only for the integer $i_{m}\in N^{+}$ such that $m\in\left(i_{m}-\frac{\Delta m}{2},i_{m}+\frac{\Delta m}{2}\right)$, so the final expression for $f_{M}(m)$ is
$f_{M}(m)$ $=\begin{cases}\frac{1-e^{-\beta\Delta m}}{\Delta m}\,e^{-\beta\Delta m\,(i_{m}-1)},&\text{if}\quad m\in\left(i_{m}-\frac{\Delta m}{2},i_{m}+\frac{\Delta m}{2}\right),\quad i_{M}\in N^{+}\\ 0&\text{otherwise},\end{cases}$ (10)
@@
We can conclude that, if $M_{i}$ is geometrically distributed with parameter $1-e^{-\beta\Delta m}$, while $Y$ is uniformly distributed in $(-\frac{\Delta m}{2},\frac{\Delta m}{2})$, the random variable $M=M_{i}+Y$ follows a residual lifetime distribution, i.e., a staircase distribution of the process of holding and jump times with probability mass allocated across steps, whose PDF $f_{M}(m)$:
@@ 1.2.1 Finite magnitudes’ dataset (real case-studies)
The calculations presented above are based on an infinite, discrete set of magnitudes, i.e., $i=1,2,\dots$. However, this assumption does not hold in real applications, where earthquake catalogs are finite and may vary significantly in quality and size, from statistically limited small datasets to more recent high-resolution records. In theoretical terms, this implies considering a truncated geometric distribution for $M_{i}$. Specifically,
@@
where we used $q=1-e^{-\beta\Delta m}$. This is again a piecewise constant, continuous PDF of a residual lifetime distribution, consisting of $N$ rectangular bumps of width $\Delta m$, centered at every $i=1,\dots,N$ and with exponentially decaying height (i.e., a staircase distribution). Specifically, in each interval:
@@
Figure 1: Numerical test showing how $\Delta m$ impacts the approximation quality of the exponential distribution (solid red line). The analytical solutions of the piecewise constant staircase PDF (a) and CDF (b) are calculated for the following values of $\Delta m$: 0.1, 0.2, 0.3, 0.4, 0.5. Since the maximum magnitude is fixed ($M_{max}=5.0$), here $N$ is a dependent variable: $N=(M_{max}-M_{0})/\Delta_{M}$, where $M_{0}=1.0$ is the minimum magnitude. Note: the staircase PDF is normalized by construction, i.e., the areas of the piecewise-constant rectangles of width $\Delta m$ sum to 1.
@@ 1.2.2 Numerical test: binned magnitudes dithered with uniform noise
In this numerical test, we aim to answer the following question: does uniform dithering bias $M_{c}$ estimates? To simulate binned magnitudes, we draw samples from the discretized GR law, i.e. the geometric distribution with parameter $p=1-e^{-\beta\Delta m}$. We generate synthetic catalogs of binned magnitudes while varying only the catalog size and keeping all other parameters fixed: bin width $\Delta m=0.1$, bb\,-value =1=1, and minimum magnitude $M_{\min}=1.0$. For each catalog, we estimate the magnitude of completeness $M_{c}$ using the Lilliefors test ($\alpha=0.1$, a conservative choice from a statistical point of view (clauset2009power)) as implemented in the Python routine Mc-Lilliefors (marcus_herrmann_2020_4162497; herrmann2021inconsistencies). In this idealized case, the true completeness magnitude is known ($M_{c}=M_{\min}=1.0$). Therefore, any systematic bias in the estimate of $M_{c}$ could be indicative of a methodological flaw.
Table 1 shows that the Lilliefors test correctly identifies $M_{c}$ (i.e., $M_{c}\sim M_{\min}=1.0$) only for catalog sizes of the order of several thousand events. For larger catalogs, the estimated $M_{c}$ becomes systematically higher, with the error increasing as the catalog size increases.
Table 1: Estimated magnitude of completeness $M_{c}$ obtained using the Lilliefors test (lilliefors1969kolmogorov) for synthetic catalogs with $M_{\min}=1.0$ and varying catalog size ($\alpha=0.1$). For each catalog size, $M_{c}$ is calculated as the mean of the estimates obtained from 50 independent synthetic catalogs. For each synthetic catalog, 100 noise realizations are considered.
@@
To examine this issue more in detail, we set up a numerical experiment to assess how the rejection probability of the Lilliefors test varies with the bin width $\Delta m$ and catalog size. For various combinations of $\Delta m$ and size, we generate 1,000 synthetic catalogs of binned magnitudes. Each catalog is then dithered following the standard procedure, i.e. by adding uniformly distributed noise. Lastly, we apply the Lilliefors test. The rejection probability is defined as the proportion of simulations in which the null hypothesis of exponentially distributed magnitudes is rejected. The full procedure is described in detail in Algorithm 1 in the Supplemental Material (Section S4). Results are shown in Table 2 and Figure 3.
Table 2: Lilliefors test rejection probability (%) for $\alpha$ = 0.1, based on 1,000 simulations with 100 noise realizations each, shown as a function of bin width ($\Delta m$) and catalog size.
@@
0.5 17.9 100 100 100 100
0.4 4.8 100 100 100 100
0.3 5.3 92.7 100 100 100
0.2 5.6 14.2 100 100 100
0.1 5.4 5.6 21.3 100 100
The rejection probability increases with both catalog size and bin width $\Delta m$. Specifically, for a fixed catalog size, the rejection probability tends to increase as $\Delta m$ increases. With the standard bin width ($\Delta m=0.1$), the exponential model is rejected 100% of the time for size of order $10^{5}$, which corresponds to the lower range of modern high‐resolution datasets (e.g. ross2019searching; tan2021machine).
@@
    For a fixed significance level, increasing catalog size and bin width leads to a higher rejection probability for the Lilliefors test, as larger samples provide more information and wider bins induce stronger deviations from the exponential model, thus increasing the power of the test.
The apparently correct estimation of $M_{c}=M_{\min}$ (Table 1) and the low rejection probability at low-to-intermediate catalog sizes (Table 2) is an artefact caused by two opposing effects compensating each other, i.e. smoothing by dithering and the statistical power of the Lilliefors test. For catalog sizes of a few thousand events, the uniform noise smooths the geometric distribution sufficiently well that the deviations from the exponential model are too small for the Lilliefors test to detect, so the null hypothesis is not rejected. As the catalog size increases, however, the test gains statistical power and begins to detect the systematic deviations introduced by the uniform dithering, causing a growing number of rejections and an increase in the estimated $M_{c}$.
Figure 3: Lilliefors test rejection probability (%) for $\alpha$ = 0.1, computed from 1,000 simulations, each incorporating 100 noise realizations, across varying bin widths ($\Delta m$) and catalog sizes. Independent random seeds are used for each bin width and catalog size combination. See Table 2 for the exact probability values.